OpenStudy (anonymous):

Linear Algebra. "*"---> means multiply Let {v1, v2,.... Vn}be an orthonormal basis for Rn, For any x and y in Rn prove that x*y = (x*v1)(v1*y) + (x*v2)(v2*y) +....+ (x*Vn)(Vn* y):

OpenStudy (amistre64):

x = x1v1 + x2v2 + x3v3+...+xnvn y = y1v1 + y2v2 + y3v3 +...+ynvn ------------------------------ might be something like this

OpenStudy (amistre64):

not quite sure what you mean by multiply tho, im reading it as a dot product tho

OpenStudy (anonymous):

yea its suppose to be a dot

OpenStudy (amistre64):

im not sure what to make of your notation for it tho; my xns and yns would correspond to the component parts that add up the basis vectors to create x and y with. what would (x*vn)(vn*y) mean?

OpenStudy (anonymous):

its question 8

OpenStudy (anonymous):

its okay if you dont get it

OpenStudy (amistre64):

id have to take it to paper to run a few ideas

OpenStudy (amistre64):

start with the orthoN , v1 = 1,0 v2 = 0,1 and try to work an example

OpenStudy (amistre64):

get a feel for it

OpenStudy (amistre64):

lets pick some random vectors for x and y x = 2,3 y = -3,8 lets use the vectors v1=1,0 and v2 = 0,1 x = <2,3> + <2,3> <1,0> <0,1> ---------------- 2 + 3 y = <-3,8> + <-3,8> <1,0> <0,1> ---------------- -3 + 8 2(-3) + 8(3) = -6 + 24 = 18 x = 2,3 y = -3,8 ----- -6+24 = 18 well, the concept works out

OpenStudy (anonymous):

yea but i think it means to prove that both sides equal each other just not sure how

OpenStudy (amistre64):

yeah, its prolly something hidden in a definitional thrm or something thats eluding us

OpenStudy (amistre64):

not sure what it gets us, but this is where i get to lol x = a y = b u = any unit vector \[|a|\cos\alpha =a\cdot u \]\[|b|\cos\beta =b\cdot u \]------------ \[|a||b|\cos\alpha \cos\beta =(a\cdot u)(b\cdot u) \]\[|a||b|\cos\gamma =a\cdot b \] the idea is that \[|a||b|\cos\alpha \cos\beta = |a||b|\cos\gamma\]  \[\cos\alpha \cos\beta = \cos\gamma\]\[\cos\alpha \cos\beta = \cos(\alpha-\beta )\] \[\cos\alpha \cos\beta = \cos\alpha\cos\beta +\sin\alpha\sin\beta\] \[0 = \sin\alpha\sin\beta;\ \alpha=0,\pi\ or\ \beta=0,\pi\]